OpenStudy (anonymous):

I am working factoriztions of 14, 40, and 42. I know the prime numbers of each but i don't know how to show my work. ex. 14= 2x7 42= 2x3x7 .

OpenStudy (anonymous):

That's all you have to do.

OpenStudy (anonymous):

There's not much work to show.

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

ok so when i check for the lcm which i believe is 1,2 is the right or a higher number

OpenStudy (anonymous):

if you wanted to be super explicit, you could show this 42= 2*21 2*21=2*3*7 42=2*3*7

OpenStudy (anonymous):

the LCM is never 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The LCM of any two numbers a and b is: \[\frac{a\cdot b}{gcd(a,b)}\]

OpenStudy (anonymous):

Did you mean the greatest common factor or greatest common denominator?

OpenStudy (anonymous):

I dont undersand that too well I ned to see numbers, please

OpenStudy (anonymous):

lcm

OpenStudy (anonymous):

Ok, so to find the LCM, you need to find the GCD

OpenStudy (anonymous):

Or you can factorize it as you have here, and just say that the LCM is the product of all the factors that the numbers have in common times the product of all of the factors that they don't have in common.

OpenStudy (anonymous):

So what numbers are you trying to find the LCM for?

OpenStudy (anonymous):

so 2 will go evenly in all the numbers 14, 40, and 42

OpenStudy (anonymous):

Ok, finding the LCM of 14, 40 and 42. 42 = 2*3*7 40 = 2*2*2*5 14 = 2*7 So yes, they all have a 2 in common. And two of them have a 7 in common. So the product of the factors they have in common is 7*2 = 14. The product of the factors they do not share is: 2*2*3*5 = 60. The product of 14 and 60 is 840

OpenStudy (anonymous):

Therefore the LCM of 14, 40 and 42 is 840

OpenStudy (anonymous):

Thanks alot you have made my day!!!